OpenStudy (anonymous):

A bag contains several marbles. Some are red some are white and some are blue. The probability of drawing a red marble is 1/6 and the probability of drawing a white marble is 1/3 . What is the probab of drawing a blue marble? What is the smallest number of marbles that could be in the bag? Could the bag contain 48 colors? If so how many of each color? If the bag contains four red marbles and eight white marbles how many blue marbles?

OpenStudy (anonymous):

For the first one I was thinking 3/6 but then I don't know and im not so sure

OpenStudy (anonymous):

1-(1/6+1/3)

Directrix (directrix):

Only red, white, and blue marbles are in the sack. Prob of red, white, or blue = 1. 1 - 1/6 - 1/3 = 3/6 or 1/2 for the prob of blue.

Directrix (directrix):

The plot thickens: What is the smallest number of marbles that could be in the bag?

OpenStudy (anonymous):

Ooh I see I just did it by doing this 1/6 + 1/3 = 3/6=1/2

OpenStudy (anonymous):

6?

Directrix (directrix):

I'm looking at the denominators of the fractional probabilities and thinking maybe the answer is the LCM of 6, 3, and 6. Least Common Multiple is 2 * 3 = 6

Directrix (directrix):

I think we should check that against the original info to see if it works.

Directrix (directrix):

Well, that is so obvious that I didn't see it. Six, yes.

Directrix (directrix):

What about this: Could the bag contain 48 colors? If so how many of each color?

OpenStudy (anonymous):

Lol and yes I thinks its correct cause 1/6 + 1/3+ 1/2= 6/6 =1

Directrix (directrix):

I can buy into that.

OpenStudy (anonymous):

I meant 48 marbles not colors

Directrix (directrix):

I'm thinking about dividing 6 into 48 and then doing something. Just an idea.

OpenStudy (anonymous):

24/48 and umm

Directrix (directrix):

Do the other two colors.

Directrix (directrix):

>>24/48 That is only the blues, right?

OpenStudy (anonymous):

No 8/48 reds and 16/48 whites

OpenStudy (anonymous):

Wait yess I guess since they add up to 24/48and thats half of 48 then theres 24/48 blue ones

Directrix (directrix):

There are two questions here: Could the bag contain 48 colors? If so how many of each color?

OpenStudy (anonymous):

I mean could the bag contain 48 marbles not colors sorry about that

Directrix (directrix):

How do you answer this one: Could the bag contain 48 marbles?

OpenStudy (anonymous):

So yes and 8/48 red 16/48 white and 24/48 blues which simplify to 1/6 1/3 and 1/2

Directrix (directrix):

Yes. 8 red, 16 white, and 24 blue I agree.

Directrix (directrix):

Now this: If the bag contains four red marbles and eight white marbles how many blue marbles?

OpenStudy (anonymous):

I put 3/4 cause I dont know their denominator so I just did this 4/48 + 8/48= 12/48 1/4 so for blue ones is 3/4 which in total 1/4+3/4 =4/4=1 ?

OpenStudy (anonymous):

Oh I just noticed that the answers we gor from before are being umm idk how to explain it watch 8/48 16/48 24/48 4/48 8/48 12/48

Directrix (directrix):

P(R) = 1/6 P(W) = 2/6 P(B) = 3/6 4 reds 4/? = 1/6 8 whites 8/? = 2/6 Here's where I am so far.

OpenStudy (anonymous):

How ? Did you get 8?

Directrix (directrix):

x = 24 4/24 + 8/24 + x/24 = 1 x will be the blues

OpenStudy (anonymous):

12/24 = 3/6

Directrix (directrix):

So, 12 blues.

Directrix (directrix):

>> How ? Did you get 8? I'm not sure which 8 you mean.

OpenStudy (anonymous):

Okay yes I understand . Oh 8 whites 8/? But I guess you divided 2 to 18/48

Directrix (directrix):

P(R) = 1/6 = 4/24 P(W) = 2/6 = 8/24 -this 8 comes from writing 2/6 as a fraction with a denominator of 24 P(B) = 3/6 = 12/24

OpenStudy (anonymous):

YeH yeah I see now from observing what you did. Thank you so much

Directrix (directrix):

You are welcome. It turned out to be a problem in writing fractions in equivalent forms with different denominators.

Directrix (directrix):

Because we are preserving the original probability fractions while writing them with the same different from original denominator.